# BEP: A Binary Error Propagation Algorithm for Binary Neural Networks Training

### v1

@@ 1 Introduction
The design of Neural Networks (NNs) with weights and activations constrained to binary values, typically $\pm 1$, is a promising direction for building models suited to resource-constrained environments. In particular, Binary Neural Networks (BNNs) offer a compelling solution for deploying Deep Learning (DL) models on edge devices and specialized hardware, where computational efficiency, power consumption, and memory footprint are critical design constraints (courbariaux2015binaryconnect; rastegari2016xnor; hubara2016binarized). Their primary advantage lies in replacing costly Floating-Point (FP) arithmetic with lightweight bitwise operations such as XNOR and Popcount, resulting in substantial reductions in computational complexity (qin2020binary; lucibello2022deep).
@@
Despite these advantages, effectively training BNNs remains challenging due to the non-differentiable nature of their activation functions. Consequently, conventional gradient-based learning algorithms cannot be directly applied, leading to two main classes of solutions. The most prevalent is Quantization Aware Training (QAT), which formulates the problem within a continuous optimization framework. In this paradigm, full-precision latent weights are maintained, and non-differentiable activations are bypassed during the backward pass using the Straight-Through Estimator (STE) (bengio2013estimating). However, reliance on real-valued computations for gradient calculation and weight updates confines the efficiency of binary arithmetic to the forward pass only (sayed2023systematic).
@@
An alternative line of research has explored purely binary, gradient-free learning rules, often inspired by principles from statistical physics (baldassi2015subdominant; cp+r). These methods operate directly on binary weights and avoid continuous surrogates. A recent extension applied this approach to binary Multi-Layer Perceptrons (MLPs) by generating local error signals at each layer using fixed random classifiers (colombo2025training). However, a fundamental limitation of this approach is that credit assignment remains local and error information does not propagate from the final output layer through the NN. This constraint makes such rules inapplicable to architectures where learning depends on end-to-end error propagation across layers, such as Recurrent Neural Networks (RNNs).
@@
From this perspective, this paper addresses the following research question: Is it possible to formulate a multi-layer, global credit assignment mechanism that back-propagates errors through the NN while operating exclusively within the binary domain? To the best of our knowledge, we introduce the first fully binary error Backpropagation (BP) algorithm capable of effectively training BNNs without relying on FP gradients. The algorithm, called Binary Error Propagation (BEP) hereafter, establishes a binary analog of the standard BP chain rule, where error signals – represented as binary vectors – are computed at the output and propagated backward through each layer of the NN. To ensure learning stability, BEP employs integer-valued hidden weights that provide synaptic inertia and mitigate catastrophic forgetting (kirkpatrick2017overcoming). Crucially, the entire forward and backward passes rely solely on efficient XNOR, Popcount, and increment/decrement operations. In summary, this work makes the following contributions:
@@
As a direct consequence, the proposed BEP algorithm eliminates the need for full-precision gradients and weight updates, enabling the exclusive use of efficient bitwise operations even during the learning phase. This drastically reduces both computational complexity and memory footprint. Experimental evaluations on multi-class classification benchmarks demonstrate test accuracy improvements of up to $6.89\%$ over the previous State-of-the-Art (SotA) algorithm (colombo2025training). Furthermore, BEP is the first solution to enable end-to-end binary training for RNN architectures, outperforming the QAT-based approach by an average of $10.57\%$ in test accuracy.
@@ 2 Related Literature
The predominant paradigm for training BNNs is QAT (courbariaux2015binaryconnect; hubara2016binarized; rastegari2016xnor). In this approach, models maintain latent full-precision parameters that are binarized during the forward pass, while gradients are computed with respect to the latent parameters using a surrogate gradient, typically a STE (bengio2013estimating). The STE approximates the derivative of the non-differentiable sign function as an identity within a bounded region, enabling the use of standard BP. Numerous subsequent works have built upon this foundation, introducing improvements such as learnable representations, enhanced architectures, and strategies to narrow the accuracy gap with full-precision models (lin2017towards; liu2020bi; adabin; schiavone2023binary). While QAT has achieved strong empirical results, it remains fundamentally a continuous optimization method applied to a discrete problem. Training relies on FP arithmetic, which prevents the full realization of BNN efficiency during learning and introduces a discrepancy between training and inference dynamics (yin2019understanding). Recent work (liu2018bireal; bulat2019xnornetpp; sarkar2024biper) has further refined QAT by incorporating residual connections and improved surrogate-gradient mechanisms to enhance gradient flow. In contrast, our approach diverges from this paradigm by eliminating the need for any real-valued parameters or surrogate gradients.
@@
A distinct line of research frames BNN training as a purely binary optimization problem. Early work in statistical physics explored combinatorial optimization techniques for training single-layer perceptrons (gardner1988space; engel2001statistical), while later studies investigated fully binary training via global heuristics such as simulated annealing (kirkpatrick1983optimization; hinton1990connectionist) and evolutionary strategies (salimans2017evolution; such2017deep; loshchilov2016cma). Although these methods avoid continuous relaxations, they explore the weight space via stochastic perturbations (e.g., random weight flips) and lack a structured, layer-wise credit assignment comparable to BP, which limits their scalability and efficiency in deep settings. Our work addresses this limitation by developing a multi-layer, gradient-free training method with a deterministic error-propagation rule.
@@
In parallel, research at the intersection of statistical physics and computational neuroscience has developed efficient local learning rules for binary neurons. Algorithms such as the Clipped Perceptron with Reinforcement (CP+R) (cp+r) and related message-passing approaches (sbpi; baldassi2015subdominant) introduce integer-valued hidden variables to represent synaptic confidence, demonstrating that single binary units can learn effectively. Multi-layer extensions of these rules generate layer-wise local error signals using fixed random classifiers, enabling training of several binary layers but still lacking end-to-end propagation of task loss (colombo2025training). This structural constraint notably precludes their application to a relevant class of recurrent sequential architectures, such as RNNs. BEP overcomes these shortcomings by introducing a global credit assignment mechanism that propagates binary error signals end-to-end, bridging the gap between fully binary optimization and multi-layer training of deep architectures.
@@ 3.1 BNN Architecture and Forward Propagation
The NN under consideration consists of two main components: a trainable binary backbone that extracts features, and a fixed task-specific output layer that maps these features to the final prediction, as shown in Figure 1(a). Given a mini-batch $\mathbf{x}=[\dots,\mathbf{x}^{\mu},\dots]$ of $bs$ input samples $\mathbf{x}^{\mu}$, where $\mu\in\{1,\dots,bs\}$, we first obtain their binary representations $\mathbf{a}_{0}^{\mu}=\textit{bin}(\mathbf{x}^{\mu})\in\{\pm 1\}^{K_{0}}$. The binarization function can be realized via median thresholding for images or thermometer encoding for tabular data (pmlr-v235-bacellar24a). The resulting binary batch $\mathbf{a}_{0}$ is then fed to the binary backbone.
Trainable BNN Backbone.
The backbone comprises a stack of $L$ fully-connected binary layers. For each layer $l\in\{1,\dots,L\}$, its state is defined by two matrices, following prior work on binary synapses (cp+r). The first is the matrix of hidden discrete weights $\mathbf{H}_{l}\in\mathbb{Z}^{K_{l}\times K_{l-1}}$. These integer-valued weights encode the synaptic inertia of each connection, providing a mechanism for stable learning and mitigating catastrophic forgetting (kirkpatrick2017overcoming). Although formulated as unbounded integers, in practice $\mathbf{H}_{l}$ is constrained to the finite range of a $B$-bit signed integer, $2^{B-1},2^{B-1}-1]$. The second is the matrix of visible binary weights $\mathbf{W}_{l}=\textit{sign}(\mathbf{H}_{l})\in\{\pm 1\}^{K_{l}\times K_{l-1}}$, representing the effective weights used during the forward pass. Here, $K_{l}$ denotes the number of neurons in layer $l$, and the sign function is applied component-wise, returning 1-1 for negative inputs and +1+1 otherwise. For each layer $l$, the pre-activations and activations are computed as $\mathbf{z}_{l}=\mathbf{W}_{l}\mathbf{a}_{l-1}$ and $\mathbf{a}_{l}=\textit{sign}(\mathbf{z}_{l})$, respectively. The output of the BNN backbone (i.e., the activations $\mathbf{a}_{L}$) is then fed to the task-specific output layer to produce the final predictions $\mathbf{\hat{y}}$.
@@ 3.3 Weight Update Mechanism
Once the target activation vector $\mathbf{a}_{l}^{*\mu}$ has been determined for layer $l$, the corresponding hidden weights $\mathbf{H}_{l}$ are updated. This update strengthens the association between the incoming activation vector $\mathbf{a}_{l-1}^{\mu}$ and the desired output pattern $\mathbf{a}_{l}^{*\mu}$. For every sample $\mu$ that triggers an update, we first compute a candidate weight-update matrix from the desired binary weights, following a procedure similar to that in prior work (sbpi; cp+r; colombo2025training). In particular, using the same strategy employed to obtain $\mathbf{a}_{l}^{*\mu}$, we maximize $\langle\mathbf{a}_{l}^{*\mu},\mathbf{W}_{l}\mathbf{a}_{l-1}^{\mu}\rangle$ with respect to $\mathbf{W}_{l}$. The solution gives the desired direction of change for the integer weights [^1] It follows from treating each row of $\mathbf{W}_{l}$ as an independent optimization and swapping $\mathbf{a}$ and $\mathbf{W}$ in Lemma 1.:
@@
Eq. 8 is the natural matrix generalization of the classical supervised Hebbian perceptron rule (rosenblatt1958perceptron) and of the Clipped Perceptron (CP) and CP+Reinforcement (CP+R) variants (sbpi; cp+r). When $K_{l}=1$, the update $\Delta\mathbf{H}_{l}^{\mu}$ reduces exactly to the CP update, which adds (or subtracts) the input vector $\mathbf{a}_{l-1}^{\mu}$ to the synaptic stability variables whenever the desired output $\mathbf{a}_{l}^{*\mu}$ is +1+1 (or 1-1). In the multi-layer setting, Eq. 8 applies this outer-product mechanism independently to each hidden neuron, with the desired activations $\mathbf{a}_{l}^{*\mu}$ serving as binary targets.
This potential update is then filtered by a binary mask $\mathbf{M}_{l}^{\mu}\in\{0,1\}^{K_{l}\times K_{l-1}}$ that selects which weights to modify. As shown in Appendix B, the resulting update is locally optimal with respect to the desired activations $\mathbf{a}_{l}^{*\mu}$. Following the sparse update strategy of (colombo2025training), the mask is constructed by partitioning the neurons of layer $l$ into subgroups and, within each group, updating only the incoming weights of the wrong perceptron deemed easiest to correct based on its pre-activation. This mechanism promotes a more uniform distribution of updates, leading to more stable training and improved generalization. If $\mathcal{M}$ denotes the set of mini-batch indices that trigger an update, we first compute each per-sample update $\Delta\mathbf{H}_{l}^{\mu}$ using the same pre-update hidden weights $\mathbf{H}_{l}$, and then apply the aggregated update as:
@@
While (colombo2025training) employs a fixed group size $\gamma$, corresponding to a constant number of updates per epoch, BEP uses a scheduled group size. Let $\Gamma_{l}=\{d\in\mathbb{N}:d\mid K_{l}\}$ denote the ordered list of positive divisors of $K_{l}$. The algorithm begins with a user-defined initial group size $\gamma_{0,l}\in\Gamma_{l}$ and increases it adaptively based on accuracy. Whenever the generalization error stagnates for a user-defined number of epochs, the group size is increased to the next larger divisor $\gamma_{t+1,l}=\min\big\{d\in\Gamma_{l}:d>\gamma_{t,l}\big\}$. If $\gamma_{t,l}$ is already the largest divisor of $K_{l}$ (the last element of $\Gamma_{l}$), it remains fixed. Because only one perceptron per group is updated, enlarging the groups gradually reduces the number of weight changes at each step, resulting in increasingly sparse updates. Empirically, this mechanism leads to a more stable optimization process in the later stages of training.
@@ 3.4 Analogy to Gradient-Based Backpropagation
The BEP procedure can be viewed as a binary reformulation of the classical BP computational graph (rumelhart1986learning). It preserves the global flow of information while substituting real-valued operations with binary, bit-wise counterparts. In standard BP, the weight update for layer $l$ follows a gradient-descent form:
@@ 4 Experimental Evaluation
In this section, we empirically evaluate the proposed BEP algorithm. The experiments have two main objectives: (i) to benchmark BEP against the SotA method for binary training of MLPs (colombo2025training), and (ii) to demonstrate the generality of BEP’s global credit assignment by applying it to binary RNNs, a setting for which local learning methods are unsuited. The Python code used in our experiments is available in a public repository [^2] The repository will be made available upon acceptance. Section 4.1 describes the experimental setup. Section 4.2 evaluates BEP on binary MLPs against both the SotA approach and QAT-based methods that rely on full-precision BP. Section 4.3 tests BEP on binary RNNs, contrasting its performance with QAT. Finally, Section 4.4 explores the effect of the gating hyperparameter $\nu$ on training dynamics.

### v2

@@ 1 Introduction
The design of Neural Networks (NNs) with weights and activations constrained to binary values, typically $\pm 1$, is a promising direction for building models suited to resource-constrained environments. In particular, Binary Neural Networks (BNNs) offer a compelling solution for deploying Deep Learning (DL) models on edge devices and specialized hardware, where computational efficiency, power consumption, and memory footprint are critical design constraints (Courbariaux et al., 2015; Rastegari et al., 2016; Hubara et al., 2016). Their primary advantage lies in replacing costly Floating-Point (FP) arithmetic with lightweight bitwise operations such as XNOR and Popcount, resulting in substantial reductions in computational complexity (Qin et al., 2020; Lucibello et al., 2022).
@@
Despite these advantages, effectively training BNNs remains challenging due to the non-differentiable nature of their activation functions. Consequently, conventional gradient-based learning algorithms cannot be directly applied, leading to two main classes of solutions. The most prevalent is Quantization Aware Training (QAT), which formulates the problem within a continuous optimization framework. In this paradigm, full-precision latent weights are maintained, and non-differentiable activations are bypassed during the backward pass using the Straight-Through Estimator (STE) (Bengio et al., 2013). However, reliance on real-valued computations for gradient calculation and weight updates confines the efficiency of binary arithmetic to the forward pass only (Sayed et al., 2023).
@@
An alternative line of research has explored purely binary, gradient-free learning rules, often inspired by principles from statistical physics (Baldassi et al., 2015; Baldassi, 2009). These methods operate directly on binary weights and avoid continuous surrogates. A recent extension applied this approach to binary Multi-Layer Perceptrons (MLPs) by generating local error signals at each layer using fixed random classifiers (Colombo et al., 2025). However, a fundamental limitation of this approach is that credit assignment remains local and error information does not propagate from the final output layer through the NN. This constraint makes such rules inapplicable to architectures where learning depends on end-to-end error propagation across layers, such as Recurrent Neural Networks (RNNs).
@@
From this perspective, this paper addresses the following research question: Is it possible to formulate a multi-layer, global credit assignment mechanism that back-propagates errors through the NN while operating exclusively within the binary domain? To the best of our knowledge, we introduce the first fully binary error Backpropagation (BP) algorithm capable of effectively training BNNs without relying on FP gradients. The algorithm, called Binary Error Propagation (BEP) hereafter, establishes a binary analog of the standard BP chain rule, where error signals – represented as binary vectors – are computed at the output and propagated backward through each layer of the NN. To ensure learning stability, BEP employs integer-valued hidden weights that provide synaptic inertia and mitigate catastrophic forgetting (Kirkpatrick et al., 2017). Crucially, the entire forward and backward passes rely solely on efficient XNOR, Popcount, and increment/decrement operations. In summary, this work makes the following contributions:
@@
As a direct consequence, the proposed BEP algorithm eliminates the need for full-precision gradients and weight updates, enabling the exclusive use of efficient bitwise operations even during the learning phase. This drastically reduces both computational complexity and memory footprint. Experimental evaluations on multi-class classification benchmarks demonstrate test accuracy improvements of up to $6.89\%$ over the previous State-of-the-Art (SotA) algorithm (Colombo et al., 2025). Furthermore, BEP is the first solution to enable end-to-end binary training for RNN architectures, outperforming the QAT-based approach by an average of $10.57\%$ in test accuracy.
@@ 2 Related Literature
The predominant paradigm for training BNNs is QAT (Courbariaux et al., 2015; Hubara et al., 2016; Rastegari et al., 2016). In this approach, models maintain latent full-precision parameters that are binarized during the forward pass, while gradients are computed with respect to the latent parameters using a surrogate gradient, typically a STE (Bengio et al., 2013). The STE approximates the derivative of the non-differentiable sign function as an identity within a bounded region, enabling the use of standard BP. Numerous subsequent works have built upon this foundation, introducing improvements such as learnable representations, enhanced architectures, and strategies to narrow the accuracy gap with full-precision models (Lin et al., 2017; Liu et al., 2020; Tu et al., 2022; Schiavone et al., 2023). While QAT has achieved strong empirical results, it remains fundamentally a continuous optimization method applied to a discrete problem. Training relies on FP arithmetic, which prevents the full realization of BNN efficiency during learning and introduces a discrepancy between training and inference dynamics (Yin et al., 2019). Recent work (Liu et al., 2018; Bulat and Tzimiropoulos, 2019; Vargas et al., 2024) has further refined QAT by incorporating residual connections and improved surrogate-gradient mechanisms to enhance gradient flow. In contrast, our approach diverges from this paradigm by eliminating the need for any real-valued parameters or surrogate gradients.
@@
A distinct line of research frames BNN training as a purely binary optimization problem. Early work in statistical physics explored combinatorial optimization techniques for training single-layer perceptrons (Gardner, 1988; Engel, 2001), while later studies investigated fully binary training via global heuristics such as simulated annealing (Kirkpatrick et al., 1983; Hinton, 1990) and evolutionary strategies (Salimans et al., 2017; Such et al., 2017; Loshchilov and Hutter, 2016). Although these methods avoid continuous relaxations, they explore the weight space via stochastic perturbations (e.g., random weight flips) and lack a structured, layer-wise credit assignment comparable to BP, which limits their scalability and efficiency in deep settings. Our work addresses this limitation by developing a multi-layer, gradient-free training method with a deterministic error-propagation rule.
@@
In parallel, research at the intersection of statistical physics and computational neuroscience has developed efficient local learning rules for binary neurons. Algorithms such as the Clipped Perceptron with Reinforcement (CP+R) (Baldassi, 2009) and related message-passing approaches (Baldassi et al., 2007; 2015) introduce integer-valued hidden variables to represent synaptic confidence, demonstrating that single binary units can learn effectively. Multi-layer extensions of these rules generate layer-wise local error signals using fixed random classifiers, enabling training of several binary layers but still lacking end-to-end propagation of task loss (Colombo et al., 2025). This structural constraint notably precludes their application to a relevant class of recurrent sequential architectures, such as RNNs. BEP overcomes these shortcomings by introducing a global credit assignment mechanism that propagates binary error signals end-to-end, bridging the gap between fully binary optimization and multi-layer training of deep architectures.
@@ 3.1 BNN Architecture and Forward Propagation
The NN under consideration consists of two main components: a trainable binary backbone that extracts features, and a fixed task-specific output layer that maps these features to the final prediction, as shown in Figure 1(a). Given a mini-batch $\mathbf{x}=[\dots,\mathbf{x}^{\mu},\dots]$ of $bs$ input samples $\mathbf{x}^{\mu}$, where $\mu\in\{1,\dots,bs\}$, we first obtain their binary representations $\mathbf{a}_{0}^{\mu}=\textit{bin}(\mathbf{x}^{\mu})\in\{\pm 1\}^{K_{0}}$. The binarization function can be realized via median thresholding for images or thermometer encoding for tabular data (Bacellar et al., 2024a). The resulting binary batch $\mathbf{a}_{0}$ is then fed to the binary backbone.
@@
Trainable BNN Backbone.
The backbone comprises a stack of $L$ fully-connected binary layers. For each layer $l\in\{1,\dots,L\}$, its state is defined by two matrices, following prior work on binary synapses (Baldassi, 2009). The first is the matrix of hidden discrete weights $\mathbf{H}_{l}\in\mathbb{Z}^{K_{l}\times K_{l-1}}$. These integer-valued weights encode the synaptic inertia of each connection, providing a mechanism for stable learning and mitigating catastrophic forgetting (Kirkpatrick et al., 2017). Although formulated as unbounded integers, in practice $\mathbf{H}_{l}$ is constrained to the finite range of a $B$-bit signed integer, $2^{B-1},2^{B-1}-1]$. The second is the matrix of visible binary weights $\mathbf{W}_{l}=\textit{sign}(\mathbf{H}_{l})\in\{\pm 1\}^{K_{l}\times K_{l-1}}$, representing the effective weights used during the forward pass. Here, $K_{l}$ denotes the number of neurons in layer $l$, and the sign function is applied component-wise, returning 1-1 for negative inputs and +1+1 otherwise. For each layer $l$, the pre-activations and activations are computed as $\mathbf{z}_{l}=\mathbf{W}_{l}\mathbf{a}_{l-1}$ and $\mathbf{a}_{l}=\textit{sign}(\mathbf{z}_{l})$, respectively. The output of the BNN backbone (i.e., the activations $\mathbf{a}_{L}$) is then fed to the task-specific output layer to produce the final predictions $\mathbf{\hat{y}}$.
@@ 3.3 Weight Update Mechanism
Once the target activation vector $\mathbf{a}_{l}^{*\mu}$ has been determined for layer $l$, the corresponding hidden weights $\mathbf{H}_{l}$ are updated. This update strengthens the association between the incoming activation vector $\mathbf{a}_{l-1}^{\mu}$ and the desired output pattern $\mathbf{a}_{l}^{*\mu}$. For every sample $\mu$ that triggers an update, we first compute a candidate weight-update matrix from the desired binary weights, following a procedure similar to that in prior work (Baldassi et al., 2007; Baldassi, 2009; Colombo et al., 2025). In particular, using the same strategy employed to obtain $\mathbf{a}_{l}^{*\mu}$, we maximize $\langle\mathbf{a}_{l}^{*\mu},\mathbf{W}_{l}\mathbf{a}_{l-1}^{\mu}\rangle$ with respect to $\mathbf{W}_{l}$. The solution gives the desired direction of change for the integer weights [^1] It follows from treating each row of $\mathbf{W}_{l}$ as an independent optimization and swapping $\mathbf{a}$ and $\mathbf{W}$ in Lemma 1.:
@@
Eq. 8 is the natural matrix generalization of the classical supervised Hebbian perceptron rule (Rosenblatt, 1958) and of the Clipped Perceptron (CP) and CP+Reinforcement (CP+R) variants (Baldassi et al., 2007; Baldassi, 2009). When $K_{l}=1$, the update $\Delta\mathbf{H}_{l}^{\mu}$ reduces exactly to the CP update, which adds (or subtracts) the input vector $\mathbf{a}_{l-1}^{\mu}$ to the synaptic stability variables whenever the desired output $\mathbf{a}_{l}^{*\mu}$ is +1+1 (or 1-1). In the multi-layer setting, Eq. 8 applies this outer-product mechanism independently to each hidden neuron, with the desired activations $\mathbf{a}_{l}^{*\mu}$ serving as binary targets.
This potential update is then filtered by a binary mask $\mathbf{M}_{l}^{\mu}\in\{0,1\}^{K_{l}\times K_{l-1}}$ that selects which weights to modify. As shown in Appendix B, the resulting update is locally optimal with respect to the desired activations $\mathbf{a}_{l}^{*\mu}$. Following the sparse update strategy of (Colombo et al., 2025), the mask is constructed by partitioning the neurons of layer $l$ into subgroups and, within each group, updating only the incoming weights of the wrong perceptron deemed easiest to correct based on its pre-activation. This mechanism promotes a more uniform distribution of updates, leading to more stable training and improved generalization. If $\mathcal{M}$ denotes the set of mini-batch indices that trigger an update, we first compute each per-sample update $\Delta\mathbf{H}_{l}^{\mu}$ using the same pre-update hidden weights $\mathbf{H}_{l}$, and then apply the aggregated update as:
@@
While (Colombo et al., 2025) employs a fixed group size $\gamma$, corresponding to a constant number of updates per epoch, BEP uses a scheduled group size. Let $\Gamma_{l}=\{d\in\mathbb{N}:d\mid K_{l}\}$ denote the ordered list of positive divisors of $K_{l}$. The algorithm begins with a user-defined initial group size $\gamma_{0,l}\in\Gamma_{l}$ and increases it adaptively based on accuracy. Whenever the generalization error stagnates for a user-defined number of epochs, the group size is increased to the next larger divisor $\gamma_{t+1,l}=\min\big\{d\in\Gamma_{l}:d>\gamma_{t,l}\big\}$. If $\gamma_{t,l}$ is already the largest divisor of $K_{l}$ (the last element of $\Gamma_{l}$), it remains fixed. Because only one perceptron per group is updated, enlarging the groups gradually reduces the number of weight changes at each step, resulting in increasingly sparse updates. Empirically, this mechanism leads to a more stable optimization process in the later stages of training.
@@ 3.4 Analogy to Gradient-Based Backpropagation
The BEP procedure can be viewed as a binary reformulation of the classical BP computational graph (Rumelhart et al., 1986). It preserves the global flow of information while substituting real-valued operations with binary, bit-wise counterparts. In standard BP, the weight update for layer $l$ follows a gradient-descent form:
@@ 4 Experimental Evaluation
In this section, we empirically evaluate the proposed BEP algorithm. The experiments have two main objectives: (i) to benchmark BEP against the SotA method for binary training of MLPs (Colombo et al., 2025), and (ii) to demonstrate the generality of BEP’s global credit assignment by applying it to binary RNNs, a setting for which local learning methods are unsuited. The Python code used in our experiments is available in a public repository [^2] https://github.com/AI-Tech-Research-Lab/BEP. Section 4.1 describes the experimental setup. Section 4.2 evaluates BEP on binary MLPs against both the SotA approach and QAT-based methods that rely on full-precision BP. Section 4.3 tests BEP on binary RNNs, contrasting its performance with QAT. Finally, Section 4.4 explores the effect of the gating hyperparameter $\nu$ on training dynamics.